OpenStudy (anonymous):

Factor completely 5a2 + b. Prime a(5a + b) b(5a2) ab(5a + 1)

OpenStudy (anonymous):

@Arianna842 @ajaja @Starr_DynastyT @SolomonZelman

OpenStudy (anonymous):

@Darknight200 @Darknight200 @Deg

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

IS THIS PRIME

OpenStudy (anonymous):

does not factor so must be prime

OpenStudy (anonymous):

the answer is prime a(5a + b) = 5a^2 + ab b(5a^2) = 5a^2b b(5a+ 1) = 5a^2b + ab

OpenStudy (anonymous):

sorry was off line

OpenStudy (anonymous):

FAN AND MEDAL AS SAID

OpenStudy (anonymous):

:p :))